OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6.

OpenStudy (anonymous):

what is this ?

Directrix (directrix):

Let x be the number x/3 - 8 = -6 This is the equation to solve for x. @aLeXiS92

Directrix (directrix):

x/3 - 8 = -6 +8 +8 ----------------- What do you get when you add 8 to both sides?